### 28 Dec 2023 v1

# Betti numbers and compactible almost complex structures on the Hard Lefschetz condition

Abstract
A special class of symplectic manifolds is those satisfying the Hard Lefschetz condition. The algebraic property of $\mathfrak{sl}_{2}(\mathbb{C})$-module gives a natural constraint on the Betti numbers. However, such a constraint is only necessary, and it is not so clear what condition on the Betti numbers is equivalent to the Hard Lefschetz condition. In this paper, we give some relations between the Betti numbers and the variant Hodge numbers on the Hard Lefschetz condition. In general, we give a sufficient condition on Betti numbers and variant Hodge numbers for any almost Kähler manifold satisfying the Hard Lefschetz condition. In particular, we offer an equivalent condition on the first Betti number for compact symplectic $4$-manifolds satisfying the Hard Lefschetz condition.
Keywords: symplectic manifolds, Hard Lefschetz condition, Dolbeault cohomology, Betti numbers 2010 Math. Subject Class. 32A25, 53C56, 53D35
1 Introduction
On a compact Kähler manifold $(M,J,\omega,g)$ with real dimension $2n$, it is known that the cup products
$[\omega^{k}\wedge]:H^{n-k}(M;\mathbb{R})\to H^{n+k}(M;\mathbb{R}),~{}k\in\{1,\ldots,n\}$ (1.1)
are isomorphic, c.f., [10, Section 6-Chapter 0]. A general compact symplectic manifold is said to satisfy the Hard Lefschetz condition, provided that the above cup products are isomorphisms. The Hard Lefschetz condition is a bridge between the Kähler manifolds and the symplectic manifolds. The isomorphisms in (1.1) provide a $\mathfrak{sl}_{2}(\mathbb{C})$-module structure on the deRham cohomology algebra of $(M,\omega)$. By the algebraic property of $\mathfrak{sl}_{2}(\mathbb{C})$-module, Thurston [16] constructed a compact symplectic $4$ manifold without Kähler structure by violating the Hard Lefschetz condition. Thurston’s work showed that the topology of symplectic manifolds is quite different to Kähler manifold, e.g., $b_{1}$ can be odd. Such a result disproved Guggenheimer’s conjecture [11]. For more symplectic manifolds with non-Kähler structure, one can refer [18]. To study the symplectic manifold, the Hard Lefschetz condition also plays an important role.
  • Mathiew [14] and Yan [19] showed that the Hard Lefschetz condition is equivalent to establish the symplectic Hodge theory, i.e., each deRham cohomology class contains a symplectic harmonic form. Recall that the notion of symplectic Hodge decomposition was given by Ehresmann and Libermann [7] and reintroduced by Brylinski [2]. Yan’s work disproved Brylinski’s conjecture.
  • Merkulov [15, Proposition 1.4] showed that the Hard Lefschetz condition is equivalent to the so called $dd^{\Lambda}$-lemma, see Theorem 2.6 for more details. Later, Tomassini and Wang [17, Theorem B] generalized this result to complex manifolds with $J$-symmetric symplectic structures.
  • Gompf and Mrowka [8] constructed symplectic $4$ manifolds satisfying the Hard Lefschetz condition but without a Kähler structure.
A compact symplectic manifolds is called a Lefschetz manifold, if it satisfies the Hard Lefschetz condition. Summarizing the above results, one has the following of inclusions:
[…]
Hence, it raises a natural question for studying compact symplectic manifolds:
Question.
What kind of compact symplectic manifolds satisfies Hard Lefschetz condition?
de Bartolomeis and Tomassini [1] gave a sufficient condition in terms of the first non-zero egienvalues of Hodge Laplacian operators for compact almost Kähler manifolds satisfying the Hard Lefschetz condition. Note that an almost Kähler manifold can be regarded as a symplectic manifold equipped with a metric. In this paper, we study the Hard Lefschetz condition in terms of Betti numbers and almost complex structures. In fact, the algebraic structure of $\mathfrak{sl}_{2}(\mathbb{C})$-representation gives the natural constraints of the Betti numbers of Lefschetz manifolds:
$b_{2k+1}\mbox{ are even},b_{k}\leq b_{k+2},~{}b_{2k}>0,\mbox{ for }0\leq k<n-1,$ (1.2)
However, one can construct a compact symplectic $4$ manifold with an even first Betti number but violating the Hard Lefschetz condition, see Example 3.1. Moreover, Gompf [9, Theorem 7.1] showed that
For a given generated group $G$ and a non-negative integer $a$, there is a compact symplectic $2n(\geq 6)$-dimensional manifolds, whose fundamental group equals to $G$ and $b_{i}\geq a$ for $22\leq i\leq n-$ 2 but violating the Hard Lefschetz condition.
Thus, (1.2) is only necessary for a symplectic manifold to become a Lefschetz one. It is natural to expect a sufficient or an equivalent condition in terms of Betti numbers for compact symplectic manifolds satisfying the Hard Lefschetz condition. By observation of (1.2), a rational guess is that such a condition should be related to Betti numbers and the symplectic form. Since each symplectic manifold can be equipped with an almost Kähler structure, we equivalently consider the relation(s) between the Betti numbers and the compatible almost complex structure on Lefschetz manifolds. The main results of this paper are as follows.
Theorem 1.1.
Let $(M,g,J,\omega)$ be a compact almost Kähler $2n$-dimensional manifold. Suppose that the equalities
$2\sum_{0\leq p\leq k}\ell^{p,2k+1-p}=b^{2k+1},~{}2\sum_{0\leq p<k}\ell^{p,2k-p}+\ell^{k,k}=b^{2k}.$ (1.3)
hold for all $k\leq n$. Then, $(M,\omega)$ satisfies the Hard Lefschetz condition.
Here, $\ell^{p,q}$’s are the generalized Hodge number on compact almost Kähler manifolds, which are defined by Cirici and Wilson in [4]. One can refer to Section 3 for their definitions. In particular, we have an equivalent condition on the first Betti number for a compact symplectic $4$-manifold to satisfy the Hard Lefschetz condition.
Theorem 1.2.
Let $(M,\omega)$ be a compact symplectic $4$ manifold. Then, $(M,\omega)$ satisfies the Hard Lefschetz condition if and only if $b_{1}=2\tilde{h}^{1,0}_{J}=\ell^{1,0}$ for any compactible almost complex structure $J$, where $\tilde{h}^{1,0}$ is the refined Hodge number.
Such a result can be used to calculate the Dolbeault cohomology on symplectic $4$ manifolds modeled on Lie groups, see Example 3.1. Furthermore, for dimension $\geq 6$ and $b_{1}\neq 0$, the Hard Lefschetz condition implies either the symplectic manifold is Kählerian or for each compactible almost complex structure its associated Nijenhuis tensor can not take the maximal rank at any point, see Proposition 3.6. The structure of this paper is as follows: In Section 2.1, we review the basic notions and properties of the Hard Lefschetz condition; In Section 2.2, we review the Dolbeault cohomology on almost complex manifold and almost Kähler manifolds and the related properties; In Section 3, we show some technique lemmas and give the proofs to the above theorems.
Acknowledgement.
The author would thank Teng Huang for introducing this topic and valuable discussion, and Adriano Tomassini for suggesting considering Example 3.1. This work is partially supported by National Natural Science Foundation of China No. 12301061 and Fundamental Research Fundsfor the Central Universities No. 2023CDJXY-043.
2 Hard Lefschetz condition and Almost complex manifolds
In this section, we review some necessary and related notions and definitions on the Hard Lefschetz condition.
2.1 Hard Lefschetz condition on linear spaces
In this subsection, we give a quick review about the Hard Lefschetz condition on linear symplectic vector space and some properties on the toy model of finite dimensional spaces, c.f., [17, Section 2 and 3]. First, we review some basic notions about the symplectic space.
Definition 2.1.
Let $V$ be a $2n$-dimensional real vector space. A bilinear form $\omega$ on $V$ is called symplectic, if $\omega$ is skew-symmetric and non-degenerate.
Similar to the metric case, the symplectic form $\omega$ defines a star operator *$**_{s}:\wedge^{p}V^{*}\rightarrow\wedge^{2n-p}V^{*}$ by
$\mu\wedge*_{s}\nu=\omega^{-1}(\mu,\nu)\frac{\omega^{n}}{n!},$
where $\omega^{-1}$ is the associated symplectic form on $V^{*}$. Set $L:\Lambda V^{*}\to\Lambda V^{*}$ as $\alpha\mapsto\omega\wedge\alpha$ and $\Lambda:=*_{s}^{-1}L*_{s}$. It is not hard to see,
$\omega^{-1}(Lu,v)=\omega^{-1}(u,\Lambda v),$
i.e., $\Lambda$ is the adjoint of $L$. For the convenience of the later arguments, we set $L^{r}:=\frac{\omega^{r}}{r!}$.
Definition 2.2.
A linear map $J:V\rightarrow V$ is called an almost complex structure on $V$ if $J^{2}=-Id$. An almost complex structure $J$ is said to be tamed by $\omega$ if
$\omega(u,Ju)>0,$
for each non-zero $u\in V$. We say $J$ is $w$-compatible if it is both taming and symmetric, i.e., $\omega(u,Jv)=\omega(v,Ju),$ for each $u,v\in V$.
An almost complex structure $J$ induces the decomposition
$V_{\mathbb{C}}:=V\otimes{\mathbb{C}}=V^{1,0}\oplus V^{0,1},$
where $V^{1,0}$ and $V^{0,1}$ are the eigenspaces of $J$ corresponding to eigenvalues $i$ and $-i$ respectively. This $J$ action can be extended to any $p$-form $\psi$ on by
$(J\psi)\left(v_{1},\cdots,v_{p}\right)=(-1)^{p}\psi\left(Jv_{1},\cdots,Jv_{p}\right).$
Similarly, we decompose $V^{*}_{\mathbb{C}}:=V^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ as $V^{*}_{\mathbb{C}}=V^{*,(1,0)}_{\mathbb{C}}\oplus V^{*,(0,1)}_{\mathbb{C}}$. Let $\bigwedge^{p,q}V^{*}$ denote the space of forms of $\bigwedge^{p,q}V^{*}:=\bigwedge^{p}V^{*,(1,0)}_{\mathbb{C}}\otimes\bigwedge^{q}V^{*,(0,1)}_{\mathbb{C}}$. An almost complex structure $J$ on $V$ compatible with $\omega$ defines an Hermitian inner product structure on $\bigwedge^{p,q}V^{*},0\leq p,q\leq n$. The associated Hodge star operator *$**:\bigwedge^{p,q}V^{*}\rightarrow\wedge^{n-q,n-p}V^{*}$ is defined by
$u\wedge*\bar{v}=(u,v)\omega^{n}/n!.$
The following identity gives a relation between *** and *$*_{s}$,
[…]
Now we give the definition of Lefschetz space and $\mathfrak{sl}_{2}(\mathbb{C})$-module, more details can be found in [[Section6-Chapter0]GH]
Definition 2.3.
Let $A=\bigoplus_{k=0}^{2n}A^{k}$ be a direct sum of complex vector spaces with possibly infinite dimensions. We say that $L\in\operatorname{End}(A)$ satisfies the Hard Lefschetz Condition and $(A,L)$ is a Lefschetz space if
$L\left(A^{l}\right)\subset A^{l+2},~{}L\left(A^{2n-1}\right)=L\left(A^{2n}\right)=0,$
for $0\leq l\leq 2(n-1)$ and each $L^{k}:A^{n-k}\rightarrow A^{n+k},0\leq k\leq n$ is an isomorphism.
We review the definitions of Lefschetz space and $\mathfrak{sl}_{2}(\mathbb{C})$-triple.
Definition 2.4.
Set $\Lambda=*_{s}^{-1}L*_{*},B:=[L,\Lambda]$. $(L,\Lambda,B)$ is called an $\mathfrak{sl}_{2}(\mathbb{C})$-triple on $(A,L)$.
The general Kähler identity implies the following result.
Theorem 2.5 (c.f. [12]).
Let $(A,L)$ be a Lefschetz space. Let $d$ be a $\mathbb{C}$-linear endomorphism of $A$ such that $d\left(A^{l}\right)\subset A^{l+1}$ and $d^{\Lambda}:=(-1)^{l+1}*_{s}d*_{s}:A^{l}\to A^{l-1}$. If $[d,L]=0$, then
  • (1)
    $(\ker\left.d\cap\operatorname{ker}d^{\Lambda},L\right)$ and $(\ker\left.d\cap\operatorname{ker}d^{\Lambda},\Lambda\right)$ are Lefschetz.
  • (2)
    $\left(\operatorname{im}d+\operatorname{im}d^{\Lambda},L\right)$ and $\left(\operatorname{im}d+\operatorname{im}d^{\Lambda},\Lambda\right)$ are Lefschetz.
  • (3)
    Assume further that $d^{2}=0$. Then, $\left(\operatorname{im}dd^{\Lambda},L\right)$ and $\left(\operatorname{im}dd^{\Lambda},\Lambda\right)$ are Lefschetz.
Let $(A,L)$ be a Lefschetz space. Let $d$ be a C-linear endomorphism of $A$ such that $d\left(A^{l}\right)\subset A^{l+1}$. We call $(A,L,d)$ a Lefschetz complex if $d^{2}=0$. Set
$H_{d}=\bigoplus_{k=0}^{2n}H_{d}^{k},\quad H_{d}^{k}:=\frac{\operatorname{ker}d\cap A^{k}}{\operatorname{Im}d\cap A^{k}},\quad H_{d^{\Lambda}}=\bigoplus_{k=0}^{2n}H_{d^{\Lambda}}^{k},\quad H_{d^{\Lambda}}^{k}:=\frac{\operatorname{ker}d^{\Lambda}\cap A^{k}}{\operatorname{Im}d^{\Lambda}\cap A^{k}}.$
One has the following theorem.
Theorem 2.6 (c.f. [17, Theorem 3.5]).
Let $(A,L,d)$ be a Lefschetz complex. Assume that $[d,L]=0$. Then the followings are equivalent:
  • (1)
    $(A,L,d)$ satisfies the $dd^{\Lambda}$-Lemma.
  • (2)
    The natural map $\ker d\cap\ker d^{\Lambda}\rightarrow H_{d}$ is surjective and the $\mathfrak{sl}_{2}(\mathbb{C})$-triple $(L,\Lambda,B)$ on $\ker d\cap\ker d^{\Lambda}$ induces the $\mathfrak{sl}_{2}(\mathbb{C})$-triple on $\mathrm{H}_{d}$.
  • (3)
    $\left(H_{d},L\right)$ satisfies the Hand Lefschetz condition.
  • (4)
    $\left(H_{d^{\Lambda}},\Lambda\right)$ satisfies the Hard Lefschetz condition.
Here a Lefschetz complex $(A,L,d)$ satisfies the $dd^{\Lambda}$-Lemma if
$\ker d\cap\ker d^{\Lambda}\cap\left(\operatorname{im}d+\operatorname{im}d^{\Lambda}\right)=\operatorname{im}dd^{\Lambda},$
on each $A^{k},0\leq k\leq 2n$.
For a compact symplectic manifold $(M,\omega)$, setting $A=\mathcal{A}^{*}$, $L=\omega\wedge$ and $d$=deRahm differential, we say $(M,\omega)$ satisfies the Hard Lefschetz condition or $(M,\omega)$ is a Lefschetz manifold, if it satisfies one of the equivalent statements in Theorem 2.6.
2.2 Cohomology groups on almost complex manifolds
In this subsection, we review the Dolbeault cohomology on almost complex manifolds and generalized Hodge number on almost Kähler manifolds. We first review some basic definitions and notions.
Definition 2.7.
For a smooth manifold $M$ of real dimension $2n$. If there exists a smooth section $J\in\Gamma(End(TM))$ on $TM$ satisfying $J^{2}=-1$, then we call $(M,J)$ an almost complex manifold with almost complex structure $J$.
Similar to the decompositions on linear toy model $V$ of the previous subsection, the parallel decompositions on $T_{\mathbb{C}}M=TM\otimes_{\mathbb{R}}\mathbb{C}$ and $T^{*}_{\mathbb{C}}M:=T^{*}M\otimes_{\mathbb{R}}\mathbb{C}$ also hold. Denoting $\mathcal{A}^{*}(M)$ by the space of sections of *$\bigwedge^{*}T^{*}_{\mathbb{C}}M$, it gives that:
$\mathcal{A}^{r}(M)=\bigoplus_{p+q=r}\mathcal{A}^{p,q}(M),$
where $\mathcal{A}^{p,q}(M)$ denotes the space of forms of $\bigwedge^{p,q}T^{*}_{\mathbb{C}}M:=\bigwedge^{p}T^{*,(1,0)}_{\mathbb{C}}M\otimes\bigwedge^{q}T^{*,(0,1)}_{\mathbb{C}}M$. By the formula
$d\psi\in\mathcal{A}^{p+2,q-1}(M)+\mathcal{A}^{p+1,q}(M)+\mathcal{A}^{p,q+1}(M)+\mathcal{A}^{p-1,q+2}(M),$
for any $\psi\in\mathcal{A}^{p,q}(M)$, the deRham differential can be decomposed as follows:
$d=\mu+\partial+\bar{\mu}+\bar{\partial}\mbox{ on }\mathcal{A}^{p,q},$
where each component is a derivation. The bi-degrees of the four components are given by
$|\mu|=(2,-1),~{}|\partial|=(1,0),~{}|\bar{\partial}|=(0,1),~{}|\bar{\mu}|=(-1,2).$
Observe that the operators $\partial,~{}\bar{\partial}$ are of the first order and the operators $\mu,~{}\bar{\mu}$ are of the zero order. Similar to the case of Riemannian manifolds, we introduce the definitions of almost Hermitian manifold and almost Kähler manifold.
Definition 2.8.
A metric $g$ on an almost complex manifold $(M,J)$ is called Hermitian, if $g$ is $J$-invariant, i.e., $g(J-,J-)=g(-,-)$. The imaginary component $\omega$ of $g$ defined by $\omega=g(J-,-)$, is a real non-degenerate $(1,1)$-form. The quadruple $(M,J,g,\omega)$ is called an almost Hermitian manifold. Moreover, when $\omega$ is $d$-closed, the quadruple $(M,J,g,\omega)$ is called an almost Kähler manifold.
Recall that by doing polar decomposition, one can equip an almost Kähler structure on any symplectic manifold. The Hermitian metric $g$ induces a unique $\mathbb{C}$-linear Hodge-*** operator
*$*:\Lambda^{p,q}\rightarrow\Lambda^{n-q,n-p},$
$\mbox{defined by }g\left(\phi_{1},\phi_{2}\right)dV=\phi_{1}\wedge*\overline{\phi_{2}}$, where $dV$ is the volume form $\frac{\omega^{n}}{n!}$ and $\phi_{1},\phi_{2}\in\Lambda^{p,q}$. For any two forms $\varphi_{1}\mbox{ and }\varphi_{2}$, we define their product by,
$(\varphi_{1},\varphi_{2})=\int_{M}\varphi_{1}\wedge*\overline{\varphi_{2}}.$
Let $\delta^{*}$ denote the formal-$L^{2}$-adjoint of $\delta$ for $\delta\in\{\mu,\partial,\bar{\partial},\bar{\mu},d\}$ with respect to the metric $g$. By the integration by part, one has the identities
$\delta^{*}=-*\bar{\delta}*\mbox{ for }\delta\in\{\mu,\partial,\bar{\partial},\bar{\mu},d\}.$
By the identity $\Delta_{\bar{\partial}}+\Delta_{\mu}=\Delta_{\partial}+\Delta_{\bar{\mu}}$ on almost Kähler manifold, c.f. [1, Lemma 3.6], one has the following theorem.
Theorem 2.9 (Cirici and Wilson [4, Theorem 4.1]).
On any compact almost Kähler manifold with dimension $2n$, it holds
$\mathcal{H}_{d}^{p,q}=\mathcal{H}_{\bar{\partial}}^{p,q}\cap\mathcal{H}_{\mu}^{p,q}=\mathcal{H}_{\partial}^{p,q}\cap\mathcal{H}_{\bar{\mu}}^{p,q}=\ker(\Delta_{\bar{\partial}}+\Delta_{\mu}),$
for all $(p,q)$. Moreover, we also have the following identities:
  • $(1)$
    (Complex conjugation)
    ¯$\mathcal{H}_{\bar{\partial}}^{p,q}\cap\mathcal{H}_{\mu}^{p,q}=\mathcal{H}_{\bar{\partial}}^{q,p}\cap\mathcal{H}_{{\mu}}^{q,p}.$
  • $(2)$
    (Hodge duality)
    *:¯$*:\mathcal{H}_{\bar{\partial}}^{p,q}\cap\mathcal{H}_{\mu}^{p,q}\rightarrow\mathcal{H}_{\bar{\partial}}^{n-q,n-p}\cap\mathcal{H}_{\mu}^{n-q,n-p}.$
  • $(3)$
    (Serre duality)
    ¯$\mathcal{H}_{\bar{\partial}}^{p,q}\cap\mathcal{H}_{\mu}^{p,q}\cong\mathcal{H}_{\bar{\partial}}^{n-p,n-q}\cap\mathcal{H}_{\mu}^{n-p,n-q}.$
For the convenience of later arguments in the next section, we set $\ell^{p,q}=\dim_{\mathbb{C}}\mathcal{H}^{p,q}_{\bar{\partial}}\cap\mathcal{H}^{p,q}_{\mu}$. One has the following constrains on Betti numbers for compact almost Kähler manifolds.
Theorem 2.10 (Cirici and Wilson [4, Theorem 4.3]).
For any compact almost Kähler manifold of dimension $2n$, the following are satisfied:
  • (1)
    In odd degrees, it establishes
    $\sum_{p+q=2k+1}\ell^{p,q}=2\sum_{0\leq p\leq k}\ell^{p,2k+1-p}\leq b^{2k+1}.$
  • (2)
    In even degrees, it holds that
    $\sum_{p+q=2k}\ell^{p,q}=2\sum_{0\leq p<k}\ell^{p,2k-p}+\ell^{k,k}\leq b^{2k},$
    with $\ell^{k,k}\geq 1$ for all $k\leq n$.
On the other hand, expanding the relation $d^{2}=0$ implies the following identities:
$\bar{\mu}\bar{\partial}+\bar{\partial}\bar{\mu}=0,~{}\bar{\mu}^{2}=0,\mbox{ and }\bar{\partial}^{2}|_{\ker(\bar{\mu})}\equiv 0(\bmod~{}im(\bar{\mu})).$
These identities induce a Dolbeault-type cohomology.
Definition 2.11 (Cirici and Wilson [5, Definition 3.1]).
The Dolbeault cohomology of an almost complex $2n$-dimensional manifold $(M,J)$ is given by
$H^{p,q}_{Dol}=H^{q}(H^{p,*}_{\bar{\mu}},\bar{\partial})=\frac{\ker(\bar{\partial}:H^{p,q}_{\bar{\mu}}\to H^{p,q+1}_{\bar{\mu}})}{im(\bar{\partial}:H^{p,q-1}_{\bar{\mu}}\to H^{p,q}_{\bar{\mu}})},$
where $H^{p,q}_{\bar{\mu}}=\frac{\ker(\bar{\mu}:\mathcal{A}^{p,q}\to\mathcal{A}^{p-1,q+2})}{im(\bar{\mu}:\mathcal{A}^{p+1,q-2}\to\mathcal{A}^{p,q})}$.
Cirici and Wilson gave the definition of the Hodge-type filtration.
On the other hand, the author gave another Dolbeault cohomology on almost complex manifolds.
Definition 2.12 (c.f. [13, Definition 3.8]).
On an almost complex manifold $(M,J)$, we define the subspace $\mathcal{A}^{*,*}_{Dol}$ of $\mathcal{A}^{*,*}$ by
$\mathcal{A}^{*,*}_{Dol}=\ker(\mu)\cap\ker(\bar{\mu})\cap\ker((\bar{\partial}+\mu)\bar{\partial})\cap\mathcal{A}^{*,*}.$
The refined Dolbeault cohomology $\tilde{H}^{*,*}_{Dol}$ is defined by
$\tilde{H}^{*,*}_{Dol}=\frac{\ker(\bar{\partial})\cap\mathcal{A}^{*,*}_{Dol}}{\bar{\partial}(\mathcal{A}^{*,*-1}_{Dol})}.$
By observation, the author showed $\ell^{1,0}=h^{1,0}=\tilde{h}^{1,0}$ on compact almost Kähler $4$ manifold and $\tilde{h}^{p,0}=h^{p,0}\leq\ell^{p,0}$, for $0\leq p\leq n$ on general compact almost Hermitian $2n$-manifolds, c.f., [13, Proposition 3.5 and Lemma 4.11]. Moreover, on compact almost complex $4$ manifolds with non-trivial almost complex structure, it holds that $\tilde{h}^{0,1}<h^{0,1}=\infty$, c.f. [13, Proposition 3.9] and [3]. We end this section by showing that the Hodge number $h^{1,0}$ is equal to $\ell^{1,0}$ on any compact almost Kähler manifold.
Lemma 2.13.
Let $(M,g,J,\omega)$ be a compact almost Kähler manifold. Then, it holds that
$h^{1,0}=\tilde{h}^{1,0}=\ell^{1,0}.$
Proof By the identities $\Delta_{\bar{\partial}}+\Delta_{\mu}=\Delta_{\partial}+\Delta_{\bar{\mu}}$ and $\Delta_{\mu}\mid_{\mathcal{A}^{1,0}}=0$, any $u\in\tilde{H}^{1,0}$ implies $u\in\ker(\Delta_{\partial}+\Delta_{\bar{\mu}})=\ker(\Delta_{\partial})\cap\ker(\Delta_{\bar{\mu}})$.
Conversely, any $u\in\mathcal{H}^{1,0}_{\bar{\partial},\mu}=\mathcal{H}^{1,0}_{\partial,\bar{\mu}}$ induces $u\in\ker(\bar{\partial})\cap\ker(\bar{\mu})$ by the same arguments.
3 Proofs
In this section, we show the proofs to the main theorems. First, we review some basic properties of the Laplacian operator. Since $\Delta_{d}$ is an elliptic operator, it is well-known that $\Delta_{d}$ is Fredholm. Hence, $\Delta_{d}$ admits a Green-function $G$ on a compact manifold. Before moving on, we give two elementary lemmas.
Lemma 3.1.
Let $(M,g)$ be a compact Riemannian manifold. Under the above notations, it holds that
$[G,*]=0,~{}[G,d]=0,~{}[G,d^{*}].$
Proof By the decomposition $\mathcal{A}^{k}=\mathcal{H}^{k}_{d}\oplus im(\Delta_{d})$, it suffices to check the three formulas for the forms in $\mathcal{H}^{k}_{d}$ and $im(\Delta_{d})$ separately. Here, we just show the first one, since the other two are similar. Recall that for any $\phi\in\mathcal{H}^{k}_{d}$, its Hodge dual *$*\phi$ is a harmonic $(\dim_{\mathbb{R}}M-k)$-form, thus
$G*\phi=*G\phi=0.$
For a $k$-form $\psi=\Delta_{d}\alpha$ with $\alpha\perp\mathcal{H}^{k}_{d}$, it holds that
$G*\psi$ $=$ $G*\Delta_{d}\alpha$
$=$ $G\Delta_{d}*\alpha$
$=$ *$*\alpha=*G\Delta_{d}\alpha=*G\psi,$
where we use the identity $[\Delta_{d},*]=0$ on the second equality. We find a form $\alpha\in(\mathcal{H}^{p,q}_{\bar{\partial}})^{\perp}$ such that $\Delta_{\bar{\partial}}\alpha=\phi$. One has that
$G\bar{\partial}\phi=G\bar{\partial}(\bar{\partial}\bar{\partial}^{*}+\bar{\partial}^{*}\bar{\partial})\alpha=G\bar{\partial}^{2}\bar{\partial}^{*}\alpha+G\bar{\partial}\bar{\partial}^{*}\bar{\partial}\alpha=G(\bar{\partial}^{2}\bar{\partial}^{*}-\bar{\partial}^{*}\bar{\partial}^{2})\alpha+\bar{\partial}\alpha=\bar{\partial}G\phi-G(\bar{\partial}^{*}\bar{\partial}^{2}f)$
$\Delta_{d}|_{\mathcal{A}^{0,1}}=(\partial+\bar{\partial})(\bar{\partial}^{*})+(\bar{\partial}^{*}+\partial^{*}+\mu^{*}+\bar{\mu}^{*})(\mu+\bar{\partial}+\partial)$
=¯¯*+¯*+¯*¯+¯*+*$=\bar{\partial}\bar{\partial}^{*}+\partial\bar{\partial}^{*}+\bar{\partial}^{*}\bar{\partial}+\bar{\partial}^{*}\partial+\partial^{*}\mu+\partial^{*}\partial+\mu^{*}\mu+\bar{\mu}^{*}\bar{\partial}$
$=\Delta_{\bar{\partial}}+\Delta_{\mu}+\Delta_{\partial}+(\partial\bar{\partial}^{*}+\bar{\partial}^{*}\partial+\bar{\mu}^{*}\bar{\partial})$
Lemma 3.2.
Let $(V,\omega)$ be a $2n$-dimensional symplectic vector space, and $J$ be a compactible (almost)complex structure on $(V,\omega)$. Then, for any $\alpha\in V^{*}$, the identity
*$*(J\alpha)=\alpha\wedge\frac{\omega^{n-1}}{(n-1)!},$
where *** is the Hodge operator with respect to the metric $\omega(-,J-)$.
Proof By the linear model of the previous section, we can find a unitary basis $\{e_{1},e_{2},\cdots e_{2n}\}$ of $V$ such that
$\omega=e^{1}\wedge e^{2}+e^{3}\wedge e^{4}+\cdots+e^{2n-1}\wedge e^{2n},$
and $J(e_{2i-1})=e_{2i},~{}J(e_{2i})=-e_{2i-1}$, where $e^{i}$ is dual to $e_{i}$ for each $1\leq i\leq n$. It suffices to check the identity of the cases $\alpha=e^{2j-1}$ and $\alpha=e^{2j}$. The formulas *$*J(e^{2j-1})=*e^{2j}=-e^{1}\wedge\cdots\widehat{e^{2j}}\wedge e^{2n}=e^{2j-1}\wedge\frac{\omega^{n-1}}{(n-1)!}$ and *$*J(e^{2j})=-*e^{2j-1}=e^{1}\wedge\cdots\widehat{e^{2j-1}}\wedge e^{2n}=e^{2j}\wedge\frac{\omega^{n-1}}{(n-1)!}$ yield the identity of the lemma.
Recall that $L^{p}=\frac{\omega^{p}}{p!}:H^{n-p}(M;\mathbb{R})\to H^{n+p}(M;\mathbb{R})$. Before showing Theorem 1.1 and 1.2, we also need the following proposition.
Proposition 3.3.
Let $(M,g,J,\omega)$ be a $2n$-dimensional compact almost Kähler manifold. Then the equality $b_{1}=2\tilde{h}^{1,0}=2\ell^{1,0}$ is equivalent to the isomorphism of the map
$L^{n-1}:H^{1}(M;\mathbb{R})\to H^{2n-1}(M;\mathbb{R}).$ (3.1)
[…]
$d^{*}(L^{n-1}\alpha)=*d(J\alpha)=*id(u-\bar{u})=0.$
Conversely, assume that the isomorphism (3.1) holds. Recall that any real-valued harmonic $1$-form $\alpha$ can be uniquely written as $\alpha=u+\bar{u}$ with $u\in\mathcal{A}^{1,0}$. Moreover, such $u$ belongs to ¯,$\mathcal{H}^{1,0}_{\bar{\partial},\mu}=\mathcal{H}^{1,0}_{d}$ if and only if $J\alpha$ is also $\Delta_{d}$-harmonic. We prove the equality by contradiction. If the equality $b_{1}=2h^{1,0}$ does not hold, then by applying Lemma 3.4 below to $\mathcal{A}^{1}_{\mathbb{R}}=\mathcal{H}^{1}_{d}\oplus(\mathcal{H}^{1}_{d})^{\perp}$, there is a $\Delta_{d}$-harmonic $1$-form still denoted by $\alpha$, such that $\alpha=u+\bar{u}$, with
$J\alpha=iu-i\bar{u}\perp\ker(\Delta_{d})\cap\mathcal{A}^{1},$
and the harmonic part of $L^{n-1}\alpha$ is nonzero. The Hodge-decomposition shows
$L^{n-1}\alpha=(L^{n-1}\alpha)_{H}+G\Delta_{d}(L^{n-1}\alpha).$
Here, $(L^{n-1}\alpha)_{H}$ denotes the harmonic part of $L^{n-1}\alpha$, which is nonzero by the isomorphism of (3.1). On the other hand, we calculate
$G\Delta_{d}(L^{n-1}\alpha)$ $=$ $G\Delta_{d}(*J\alpha)$
$=$ *$*G\Delta_{d}(J\alpha)$
$=$ *$*G\Delta_{d}(iu-i\bar{u})=*(iu-i\bar{u}),$
where we use Lemma 3.1 for the second equality and $(u-\bar{u})\perp\mathcal{H}^{1}_{d}$ for the last one. Hence, the above formula gives
$\|\alpha\|^{2}_{L^{2}}=\|J\alpha\|^{2}_{L^{2}}=\|*J\alpha\|^{2}_{L^{2}}=\|L^{n-1}\alpha\|^{2}_{L^{2}}=\|G\Delta_{d}(L^{n-1}\alpha)\|^{2}_{L^{2}}.$
It implies $(L^{n-1}\alpha)_{H}=0$, contradicting to the assumption on $u$. Therefore, we proved the proposition.
Lemma 3.4.
Let $V$ be a possibly infinite dimensional real vector space with a complex structure $J$ and a $J$-compactible metric $g$. Suppose that there is a decomposition $V=V_{1}\oplus V_{2}$ with $V_{1}\perp V_{2}$ with respect to $g$ and $\dim V_{1}<\infty$. Then, $J$ preserves $V_{1}$, if and only if there is no non-trivial subspace $V^{\prime}_{1}\subset V_{1}$ such that $J(V^{\prime}_{1})\subset V_{2}$.
Proof When $J$ preserves $V_{1}$, it is clear that there is no non-zero element in $V_{1}$ which maps onto $V_{2}$. Here, we only show the “if” part, by using the proof of contradiction.
Assume that $J$ does not preserve $V_{1}$. Then, the intersection $W:=J(V_{1})\cap V_{1}$ is a proper subspace of $V_{1}$. Since $V_{1}$ is of finite dimension, one can find an orthonormal basis $\{e_{j}\}_{1\leq j\leq\dim V_{1}}$ of $V_{1}$ and a maximal integer $n$ such that
$J(e_{2i-1})=e_{2i},~{}J(e_{2i})=-e_{2i-1},\mbox{ for }1\leq i\leq n,\mbox{ and}W=\mbox{span}\{e_{j}\}_{1\leq j\leq 2n}$
and
$J(e_{s})\not\in V_{1},\mbox{for }2n<s\leq\dim V_{1}.$
If $J(e_{s})$ has a non-zero $V_{1}$ component, then $J(e_{s})\cap V_{1}(\neq 0)\subset W$. This violates the choice of $n$. Thus, $V^{\prime}_{1}:=\mbox{span}\{e_{s}\}_{2n<s\leq N}$ is a proper subspace of $V_{1}$ such that $J(V^{\prime}_{1})\subset V_{2}$, which is a contradiction.
Now, we give a sufficient condition for compact symplectic manifolds satisfying the Hard Lefschetz condition in terms of Betti numbers and $\ell^{p,q}$s.
Proof of Theorem 1.1 Since the two equalities hold, each $12k+$ 1-degree real harmonic form $\alpha_{2k+1}$ with $10\leq k<n-$ 1 can be expressed as
$\alpha_{2k+1}=\sum_{p=0}^{k}u_{p,2k+1-p}+\bar{u}_{p,2k+1-p},$
where $u_{p,2k+1-p}\in\mathcal{H}^{p,2k+1}_{\bar{\partial},\mu}=\mathcal{H}^{p,2k+1-p}_{d}$, and each $2k$-degree real harmonic form $\alpha_{2k}$ can be expressed as
$\alpha_{2k}=\sum_{0\leq p\leq k}u_{p,2k-p}+\bar{u}_{p,2k-p}+u_{2k}$
By the formula $[d^{*},L]=d^{c}$ in [1, Formula (15)], where $d^{c}=J^{-1}dJ$, it yields that
$dd^{*}(L^{n-k-1}\alpha_{2k+1})$ $=$ $(n-k-1)(L^{n-k-2}dd^{c}\alpha_{2k+1})$
$=$ $(n-k-1)\frac{\omega^{n-k-2}}{(n-k-2)!}dJ^{-1}d\sum_{0\leq p\leq k}(i^{2p-2k-1}u_{p,2k+1-p}-i^{2p-2k-1}\bar{u}_{p,2k+1-p})$
$=$ $0$
Thus, the term $L^{n-2k-1}\alpha$ is a real harmonic form, and the map
$L^{n-k-1}:H^{2k+1}_{dR}(M;\mathbb{R})\to H^{2n-2k-1}_{dR}(M;\mathbb{R}),$
is an isomorphism. By the similar arguments, the map
$L^{n-k}:H^{2k}_{dR}(M;\mathbb{R})\to H^{2n-2k}_{dR}(M;\mathbb{R}),$
is also an isomorphism for each $k$.
Proof of Theorem 1.2 For any compactible almost complex structure $J$, Proposition 3.3 and Lemma 2.13 show that the Hard Lefschetz condition is equivalent to the relation $b_{1}=2\tilde{h}^{1,0}=2\ell^{1,0}$, which finishes the proof.
The following example shows that the even condition on $b_{1}$ is not sufficient to make compact symplectic manifolds satisfy the Hard Lefschetz condition.
Example 3.1.
[…]
$[e_{1},e_{2}]=e_{3},~{}[e_{1},e_{3}]=e_{4}.$
Then, on the dual algebra $\mathfrak{g}^{*}$, it holds that
$de^{1}=0,de^{2}=0,de^{3}=-e^{1}\wedge e^{2},de^{4}=-e^{1}\wedge e^{3}.$
Recall that the Lie group $N$ admits a lattice $\Gamma$. Let $M=\Gamma\backslash N$ and the form $\omega=e^{1}\wedge e^{4}+e^{2}\wedge e^{3}$ be a symplectic structure. It is clear that $b_{1}(M)=4$. But $(M,\omega)$ does not satisfy the Hard Lefschetz condition, since
$[\omega\wedge e^{1}]=[e^{1}\wedge e^{2}\wedge e^{3}]=[d(e^{4}\wedge e^{2})]=0.$
Moreover, we also have $\tilde{h}^{1,0}=1$.
We end this paper by showing the following proposition for higher dimensional symplectic satisfying the Hard Lefschetz condition. First, we give a technique lemma.
Lemma 3.5.
Let $(M,J)$ be a compact almost complex $2n$-manifold $n\geq 3$. If the Nijenhuis tensor takes the maximal rank at some point of $M$, then $H^{1,0}_{Dol}=\tilde{H}^{1,0}_{Dol}=\{0\}$. Moreover, if the Nijenhuis tensor is nowhere vanishing, then $H^{1,0}_{Dol}=\tilde{H}^{1,0}_{Dol}=\tilde{H}^{0,1}_{Dol}=\{0\}$.
Proof An easy calculus yields
$rank_{\mathbb{C}}\mathcal{A}^{1,0}=rank_{\mathbb{C}}\mathcal{A}^{0,1}=n,~{}rank_{\mathbb{C}}\mathcal{A}^{2,0}=rank_{\mathbb{C}}\mathcal{A}^{0,2}=\frac{n(n-1)}{2}.$
For an Hermitian metric $g$, it is clear that ¯*$\bar{\partial}^{*}u=0$ and $u\in\ker(\bar{\partial})$ is equivalent to $u\in\ker(\Delta_{\bar{\partial}})$ for any $u\in\mathcal{A}^{1,0}$. Hence, all elements in $\tilde{H}^{1,0}$ are $\Delta_{\bar{\partial}}$-harmonic. Since the maximal rank condition is open, there exists a small open subset $U$ such that the restriction $N|_{U}$ is also of the maximal rank. When $n\geq 3$, it is clear that $\bar{\mu}|_{U}:\mathcal{A}^{1,0}(U)\to\mathcal{A}^{0,2}(U)$ is either isomorphic($n=3$) or injective($n>3$) by the algebraic calculus. The equation $\bar{\mu}(u)=0$ gives that $u|_{U}=0$ for any $u\in\mathcal{A}^{1,0}$. Hence, the unique-continuity-property for harmonic forms shows that $u=0$ for any $u\in\tilde{H}^{1,0}_{Dol}$.
Similarly, when $N$ is nowhere vanishing, $\mu:\mathcal{A}^{0,1}\to\mathcal{A}^{2,0}$ is either isomorphic($n=3$) or injective($n>3$). Therefore, $\mu(v)=0$ implies $v=0$ for any $v\in\mathcal{A}^{0,1}$, which proves the last statement.
Proposition 3.6.
Let $(M,\omega)$ be a compact symplectic manifold with dimension $\geq 6$. If $b_{1}\neq 0$ and satisfies the Hard Lefschetz condition, then either $(M,\omega)$ is Kählerian or for each compactible almost complex structure its associated Nijenhuis tensor can not take the maximal rank at any point.
Proof We argue by contradiction. Assume that $(M,\omega)$ is non-Kählerian. For any $\omega$-compactible almost complex structure $J$, $b_{1}=2\tilde{h}^{1,0}$ by Proposition 3.3.
If the associated Nijenhuis tensor $N$ takes maximal rank at a point $p$, then there exists a small neighborhood $U_{p}$ such that $N\mid_{U_{p}}$ is of maximal rank by the open condition of the maximal rank. Again, by the relation $\bar{\mu}|_{U_{p}}:\mathcal{A}^{1,0}(U)\to\mathcal{A}^{0,2}(U)$ is either isomorphic($n=3$) or injective($n>3$), any $u\in\tilde{H}^{1,0}_{Dol}$ is zero inside $U_{p}$. Thus, $u\equiv 0$ by the unique continuity property, which contradicts to the assumption $b_{1}\neq 0$.